# Large-Scale UWB Anchor Calibration and One-Shot Localization Using Gaussian Process

### v1

@@ Abstract
Ultra-wideband (UWB) is gaining popularity with devices like AirTags for precise home item localization but faces significant challenges when scaled to large environments like seaports. The main challenges are calibration and localization in obstructed conditions, which are common in logistics environments. Traditional calibration methods, dependent on line-of-sight (LoS), are slow, costly, and unreliable in seaports and warehouses, making large-scale localization a significant pain point in the industry. To overcome these challenges, we propose a UWB-LiDAR fusion-based calibration and one-shot localization framework. Our method uses Gaussian Processes to estimate anchor position from continuous-time LiDAR Inertial Odometry with sampled UWB ranges. This approach ensures accurate and reliable calibration with just one round of sampling in large-scale areas, I.e., 600x450 m². With the LoS issues, UWB-only localization can be problematic, even when anchor positions are known. We demonstrate that by applying a UWB-range filter, the search range for LiDAR loop closure descriptors is significantly reduced, improving both accuracy and speed. This concept can be applied to other loop closure detection methods, enabling cost-effective localization in large-scale warehouses and seaports. It significantly improves precision in challenging environments where UWB-only and LiDAR-Inertial methods fall short, as shown in the video https://youtu.be/oY8jQKdM7lU. We will open-source our datasets and calibration codes for community use.
@@ I Introduction
Ultra-Wideband (UWB) [1] is a short-range, energy-efficient radio communication technology mainly used for precise location detection and relative range measurement [2]. Devices such as Apple AirTags and Android SmartTags are gaining popularity, with UWB technology enabling them to track and locate household items accurately.
While it has seen success in applications like sports, such as football tracking, UWB still faces significant challenges in large-scale real-time industrial localization applications [3]. Most UWB studies are confined to smaller or indoor environments operating in absolute positioning (AP) mode [1, 4], as shown in Fig. 1. This condition, all anchors must be calibrated [5], and the tag requires constant range measurements from multiple anchors, which becomes impractical in occluded environments. This limitation reduces the practicality of UWB in massive areas like seaports or warehouses, where obstructions present significant hurdles.
@@
Adopting UWB in large outdoor environments presents challenges due to frequent non-line-of-sight (NLoS) issues and interference from other devices, which limit the effective measurement range. In the worst case, there could be zero connection between anchors because of the newly added container, as shown in Fig. 1. The absence of reliable GPS signals further complicates calibration. Even with calibration, the system struggles to localize in NLoS conditions, creating a significant bottleneck for field robotics and logistics automation.
In this work, we propose a comprehensive approach to address the challenges of large-scale UWB calibration and one-shot localization. Our method leverages continuous-time LiDAR-Inertial Odometry (CT-LIO) to generate a trajectory that can be sampled at any point. By interpolating this trajectory, we align each UWB measurement with its corresponding position. We then iteratively refine anchor position estimates using a Gaussian Process. Despite calibration, the system remains unable to localize in NLoS mode using standard UWB techniques [6]. To overcome this, we combine our solution with a fast descriptor-based method [7] to enhance one-shot global localization in large and repetitive environments. Our approach significantly improves existing localization methods and introduces new possibilities previously unexplored. Our contributions can be summarized as follows:
@@
    The paper introduces a method combining UWB data with CT-LIO and Gaussian Processes to accurately calibrate UWB anchors in large environments using just one round of sampling, enhancing localization in obstructed and repetitive environments, where UWB or LiDAR alone would fail.
@@ II Related Work
Figure 2: The system leverages CT-LIO to generate UWB samples for non-parametric Gaussian process fitting, calibrating anchor positions. Using these calibrated positions, a range-based search is integrated with the descriptor method for one-shot localization. Additionally, over 100 scans were collected to create a prior map for ground truth verification.
@@ III Problem Statement
Our eventual goal is to estimate the robot’s position in large-scale repetitive environments accurately. To do this, we first obtain UWB measurements $\zeta\in\mathbb{R}$ and 3D poses $\mathbf{p}(t)\in\mathbb{R}^{3}$ from a LiDAR SLAM framework, using B-spline interpolation. Then, these interpolated poses and samples are integrated into a Gaussian Process model to calibrate UWB anchor positions, which in turn improves LiDAR localization in repetitive logistic environments.
@@ III-B Interpolated UWB poses
With LiDAR running at 10Hz and UWB running at 20Hz, there are time bias and net frequency differences in the time of acquisition. To obtain the exact UWB sample pose $T(t)=(R(t),\mathbf{p}(t))$ at any given UWB sampling time $t$, we perform B-spline interpolation using the control points $T_{i}$. $R(t)$ is the rotation matrix omitted in the calibration process. To interpolate the pose $T(t)$ at a given time $t$, the process begins by identifying the knot interval $[t_{i},t_{i+1})$ such that $t$ lies within this interval. The normalized time $s$ within the interval is then computed as $s=\frac{t-t_{i}}{t_{i+1}-t_{i}}$. Next, the blending coefficients $\lambda_{j}$ are calculated from the B-spline basis matrix $B^{(N)}$ and the powers of $s$ by applying the equation $[\lambda_{0},\lambda_{1},\dots,\lambda_{N-1}]^{\top}=B^{(N)}[1,s,\dots,s^{N-1}]^{\top}$. These coefficients determine the influence of each control point on the interpolated pose.
Finally, the position component $\mathbf{p}(t)$ of the sample pose can be interpolated as:
@@ III-C Gaussian Process for UWB Anchor Calibration
Gaussian Process [17] provides a non-parametric framework for determining the underlying distribution of UWB anchors in environment $\xi\subset\mathbb{R}^{3}$. The unknown distribution is modeled as a continuous function $f(\mathbf{p})$, which represents the range to UWB anchor at a position $\mathbf{p}$, $\mathbf{p}\subset\xi$. Given a set of samples $\mathbf{p}$, generated from SLICT, we define a Gaussian Process: $f(\mathbf{p})\sim\mathcal{GP}(m(\mathbf{p}),k(\mathbf{p},\mathbf{p}^{\prime}))$ that models the relationship between these SLAM-derived poses and the UWB ranging values with set of means and covariance kernel. Here, $m(\mathbf{p})$ is the mean function, typically set to zero in the absence of prior information about the anchor positions. $k(\mathbf{p},\mathbf{p}^{\prime}))$ is the covariance function (kernel), defining the relationship between different positions $\mathbf{p}$ and $\mathbf{p}^{\prime}$.
@@
where $\nu$ and $\ell$ are the smoothness and length-scale parameters, respectively, which are adjusted based on the environmental complexity and data variability.
The predictive mean $\mu(\mathbf{p}_{*})$ and covariance $\sigma^{2}(\mathbf{p}_{*})$ of the GP can be obtained as:
$\mu(\mathbf{p}_{*})$ $=k(\mathbf{p}_{*},\mathbf{P})[k(\mathbf{P},\mathbf{P})+\sigma^{2}I]^{-1}\zeta,$ (1)
$\sigma^{2}(\mathbf{p}_{*})$ $=k(\mathbf{p}_{*},\mathbf{p}_{*})-k(\mathbf{p}_{*},\mathbf{P})[k(\mathbf{P},\mathbf{P})+\sigma^{2}I]^{-1}k(\mathbf{P},\mathbf{p}_{*}),$ (2)
@@
We use an iterative Gaussian Process to refine predictions by first fitting initial points at a coarse scale, then sampling within a rectangular cuboid region at regular intervals to account for unknown positions. The top 10 sampled values are averaged to update the model. During training, layered random sampling ensures data coverage across the region, improving generalization and convergence towards optimal results. This calibration process enables more accurate anchor positioning, which is critical for improving overall localization performance.
@@ IV-B Experiment Setup
Experiments were conducted on a desktop PC equipped with an i9-13900K CPU and an RTX 4090 GPU for the calibration phase. For the localization experiment, a Gen 11 i7 NUC was used to simulate edge industrial usage scenarios. The ground truth is captured by a map constructed by hundreds of survey scans and is accurate up to cm level. A set of Here+ RTK GPS bases with GPS receivers was deployed to collect precise GPS data, which was later integrated into the existing dataset for comparison. RTK base stations were placed strategically on the rooftop of the NTU S2 building to ensure optimal signal reception and accuracy under the given conditions. The experiment included 10 Noop LTP-A type UWB anchors and a tag covering an area of approximately 600 x 450 m². The trajectory data was primarily collected using an all-terrain vehicle (ATV) equipped with two UWB tags mounted on its body. However, only one UWB tag was used in this experiment, as the current Gaussian Process (GP) formulation does not account for the differential angle of arrival between the tags.
@@ IV-F Result and Discussion
In contrast, the proposed method, using SLICT, provided efficient and accurate pose estimates, achieving an average anchor calibration accuracy of around 2 meters. The remaining error can be attributed to odometry drift and UWB biases at longer ranges in tropical humid conditions. The only exception was Anchor 0, which was placed at B5 inside the building, while trajectory samples were collected around B6, leading to the failure of the Gaussian Process to locate its position. The overall process was efficient, as we spent 5 minutes riding the ATV to each location, followed by about 50 minutes of UWB anchor calibration process.
@@ VI Limitation and future work
Reliance on High-Performance LiDAR: Another issue is the reliance on the costly 128-line LiDAR, which limits its feasibility for large-scale operations. Most service AMRs use 16-line LiDARs or vision-based approaches. Future work will focus on improving accuracy and cost efficiency by estimating dynamic biases and replacing CT-LIO with a more affordable vision-based approach, such as AirVO [27], [28], [29], to reduce reliance on expensive LiDARs.
@@ VII Conclusion
We propose a UWB calibration method to address NLoS issues using SLICT to generate a continuous-time trajectory. By sampling UWB acquisition poses and applying an iterative Gaussian process, we successfully calibrate anchors in large 600x450 m² environments with approximately 2-meter accuracy. Even in NLoS conditions, where traditional UWB localization fails, the calibrated anchor positions serve as a stand-alone plugin to improve one-shot localization success rates in repetitive and large-scale environments. We demonstrate that container-carrying AMRs can achieve accurate one-shot localization in repetitive environments using sparsely populated UWB and onboard LiDAR at minimal cost.

### v2

@@ Abstract
Ultra-wideband (UWB) is gaining popularity with devices like AirTags for precise home item localization but faces significant challenges when scaled to large environments like seaports. The main challenges are calibration and localization under obstructed conditions, which are common in logistics environments. Traditional calibration methods, dependent on line-of-sight (LoS), are slow, costly, and unreliable in seaports and warehouses, making large-scale localization a significant pain point in the industry. To overcome these challenges, we propose a one-shot calibration and localization framework based on UWB-LiDAR fusion. Our method uses Gaussian processes to estimate the anchor position from continuous-time LiDAR Inertial Odometry with sampled UWB ranges. This approach ensures accurate and reliable calibration with only one round of sampling in large-scale areas, i.e., 600x450 $m^{2}$. With LoS issues, UWB-only localization can be problematic, even when anchor positions are known. We demonstrate that by applying a UWB-range filter, the search range for LiDAR loop closure descriptors is significantly reduced, improving both accuracy and speed. This concept can be applied to other loop closure detection methods, enabling cost-effective localization in large-scale warehouses and seaports. It significantly improves precision in challenging environments where the UWB-only and LiDAR-Inertial methods fail, as shown in the video https://youtu.be/oY8jQKdM7lU. We will open-source our datasets and calibration codes for community use.
@@
This work has been accepted to IEEE International Conference on Robotics and Automation (ICRA) © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/redistribution, creating new works, or reuse of any copyrighted components of this work in other media.
@@ I Introduction
Ultra-Wideband (UWB) [1] is a short-range, energy-efficient radio communication technology mainly used for precise location detection and relative range measurement [2]. Devices such as Apple AirTags and Android SmartTags are gaining popularity, with UWB technology that enables them to track and locate household items accurately [2].
Although it has seen success in sports-like applications, such as football tracking, UWB still faces significant challenges in large-scale real-time industrial localization applications [3]. Most UWB studies are confined to smaller or indoor environments operating in absolute positioning (AP) mode [1, 4], as shown in Fig. 1. In this setup, all anchors must be calibrated [5], and the tag requires constant range measurements from multiple anchors, which becomes impractical in occluded environments [6]. This limitation reduces the practicality of UWB in massive areas such as seaports or warehouses, where obstructions present significant hurdles.
@@
Adopting UWB in large outdoor environments [7, 8] presents challenges due to frequent Non-Line-of-Sight (NLoS) issues and interference from other devices, which limit the effective sensing range. In the worst case, there could be zero connection between the anchors due to the newly added container, as shown in Fig. 1. The absence of reliable GPS signals in obstructed environments further complicates anchor calibration and requires alternative approaches [9, 10, 11, 12, 13, 14, 15] to assist with calibration. Even with calibration [16, 17, 18], the system struggles to localize [19, 20] in NLoS conditions [21], creating a significant bottleneck for field robotics and logistics automation.
In this work, we propose a comprehensive approach to address the challenges of large-scale UWB calibration and one-shot localization. Our method leverages Continuous-Time LiDAR-Inertial Odometry (CT-LIO) to generate a trajectory that can be sampled at any point. By interpolating this trajectory, we align each UWB measurement with its corresponding position. The anchor position estimates are iteratively refined using Gaussian Process (GP) [22] regression, which models spatial dependencies and improves accuracy. We use GP for calibration instead of traditional Nonlinear Least Squares (NLS), as it allows for flexible modeling of spatially correlated UWB errors and provides a probabilistic framework for uncertainty quantification. Even after calibration, the system remains unable to localize in NLoS mode using standard UWB techniques [23]. To overcome this, we combine our solution with a fast descriptor-based method [24] to enhance one-shot global localization in large and repetitive environments. Our approach significantly improves existing localization methods and introduces new possibilities previously unexplored. Our contributions can be summarized as follows:
@@
    The paper introduces a framework that combines UWB data with CT-LIO and GP to accurately calibrate UWB anchors in large environments using only one round of sampling. This enhances localization in obstructed scenes where UWB or LiDAR alone would fail.
@@ II Related Work
Figure 2: The system leverages CT-LIO to generate UWB samples for non-parametric GP fitting, calibrating anchor positions. Using these calibrated positions, a range-based search is integrated with the descriptor method for one-shot localization. Additionally, over 100 scans were collected to create a prior map for ground truth verification.
@@ III Problem Statement
Our eventual goal is to estimate the robot’s position in large-scale repetitive environments accurately. To do this, we first obtain UWB measurements $\zeta\in\mathbb{R}$ and 3D poses $\mathbf{p}(t)\in\mathbb{R}^{3}$ from a LiDAR SLAM framework, using B-spline interpolation. Then, these interpolated poses and samples are integrated into a GP model to calibrate UWB anchor positions, which in turn improves LiDAR localization in repetitive logistic environments.
@@ III-B Interpolated UWB poses
With LiDAR running at 10Hz and UWB running at 20Hz, there are time bias and net frequency differences in the time of acquisition. To obtain the exact UWB sample pose $T(t)=(R(t),\mathbf{p}(t))$ at any given UWB sampling time $t$, we perform B-spline interpolation using the control points $T_{i}$. $R(t)$ is the rotation matrix omitted in the calibration process. To interpolate the pose $T(t)$ at a given time $t$, the process begins by identifying the knot interval $[t_{i},t_{i+1})$ such that $t$ lies within this interval. The normalized time $s$ within the interval is then computed as $s=\frac{t-t_{i}}{t_{i+1}-t_{i}}$. Next, the blending coefficients $\lambda_{j}$ are calculated from the B-spline basis matrix $B^{(N)}$ and the powers of $s$ by applying the equation $[\lambda_{0},\lambda_{1},\dots,\lambda_{N-1}]^{\top}=B^{(N)}[1,s,\dots,s^{N-1}]^{\top}$. These coefficients determine the influence of each control point on the interpolated pose. Finally, the position component $\mathbf{p}(t)$ of the sample pose can be interpolated as:
@@ III-C Gaussian Process for UWB Anchor Calibration
GP [34] provides a non-parametric framework for determining the underlying distribution of UWB anchors in environment $\xi\subset\mathbb{R}^{3}$. The unknown distribution is modeled as a continuous function $f(\mathbf{p})$, which represents the range to UWB anchor at a position $\mathbf{p}$, $\mathbf{p}\subset\xi$. Given a set of samples $\mathbf{p}$, generated from SLICT, we define a GP: $f(\mathbf{p})\sim\mathcal{GP}(m(\mathbf{p}),k(\mathbf{p},\mathbf{p}^{\prime}))$ that models the relationship between these SLAM-derived poses and the UWB ranging values with set of means and covariance kernel. Here, $m(\mathbf{p})$ is the mean function, typically set to zero in the absence of prior information about the anchor positions. $k(\mathbf{p},\mathbf{p}^{\prime}))$ is the covariance function (kernel), defining the relationship between different positions $\mathbf{p}$ and $\mathbf{p}^{\prime}$.
@@
where $\nu$ and $\ell$ are the smoothness and length-scale parameters, respectively, which are adjusted based on the environmental complexity and data variability. The predictive mean $\mu(\mathbf{p}_{*})$ and covariance $\sigma^{2}(\mathbf{p}_{*})$ of the GP can be obtained as:
$\mu(\mathbf{p}_{*})$ $=k(\mathbf{p}_{*},\mathbf{P})[k(\mathbf{P},\mathbf{P})+\sigma^{2}I]^{-1}\zeta,$
$\sigma^{2}(\mathbf{p}_{*})$ $=k(\mathbf{p}_{*},\mathbf{p}_{*})-k(\mathbf{p}_{*},\mathbf{P})[k(\mathbf{P},\mathbf{P})+\sigma^{2}I]^{-1}k(\mathbf{P},\mathbf{p}_{*}),$
@@
We use an iterative GP to refine predictions by first fitting initial points at a coarse scale, then sampling within a rectangular cuboid region at regular intervals to account for unknown positions. The top 10 sampled values are averaged to update the model. During training, layered random sampling ensures data coverage across the region, improving generalization and convergence towards optimal results. This calibration process enables more accurate anchor positioning, which is critical for improving overall localization performance.
@@ IV-B Experiment Setup
Experiments were conducted on a desktop PC equipped with an i9-13900K CPU and an RTX 4090 GPU for the calibration phase. For the localization experiment, a Gen 11 i7 NUC was used to simulate edge industrial usage scenarios. The ground truth is captured by a map constructed by hundreds of survey scans and is accurate up to cm level. A set of Here+ RTK GPS bases with GPS receivers was deployed to collect precise GPS data, which was later integrated into the existing dataset for comparison. RTK base stations were placed strategically on the rooftop of the NTU S2 building to ensure optimal signal reception and accuracy under the given conditions. The experiment included 10 Noop Loop LTP-A TOF-type UWB anchors and a tag covering an area of approximately 600 x 450 m². The trajectory data was primarily collected using an all-terrain vehicle (ATV) equipped with two UWB tags mounted on its body. However, only one UWB tag was used in this experiment, as the current GP formulation does not account for the differential angle of arrival between the tags.
@@ IV-F Result and Discussion
In contrast, the proposed method, using SLICT, provided efficient and accurate pose estimates, achieving an average anchor calibration accuracy of around 2 meters. The remaining error can be attributed to odometry drift and UWB biases at longer ranges in tropical humid conditions. The only exception was Anchor 0, which was placed at B5 inside the building, while trajectory samples were collected around B6, leading to the failure of the GP to locate its position. The overall process was efficient, as we spent 5 minutes riding the ATV to each location, followed by about 50 minutes of UWB anchor calibration process.
@@ VI Limitation and future work
Reliance on High-Performance LiDAR: Another issue is the reliance on the costly 128-line LiDAR, which limits its feasibility for large-scale operations. Most service autonomous mobile robots (AMR) use 16-line LiDARs or vision-based approaches. Future work will focus on improving accuracy and cost efficiency by estimating dynamic biases and replacing CT-LIO with a more affordable vision-based approach, such as AirVO [44], [45], [46], to reduce reliance on expensive LiDARs.
@@ VII Conclusion
We propose a UWB calibration method to address NLoS issues using SLICT to generate a continuous-time trajectory. By sampling UWB acquisition poses and applying an iterative GP, we successfully calibrated anchors in large 600x450 m² environments with approximately 2-meter accuracy. Even in NLoS conditions, where traditional UWB localization fails, the calibrated anchor positions serve as a standalone plugin to improve one-shot localization success rates in repetitive and large-scale environments. We demonstrate that container-carrying AMRs can achieve accurate one-shot localization in repetitive environments using sparsely populated UWBs and LiDAR onboard at minimal cost.